Your parents will retire in 18 years. They currently have $250,000, and they think they will need
$1 million at retirement. What annual interest rate must they earn to reach their goal, assuming
they don’t save any additional fund?

Answer by Theo(13342)

formula to use is f = p * (1 + r) ^ n
f is the future value
p is the present value
(1 + r) is the annual growth rate.
n is the number of years.

formula becomes:
1,000,000 = 250,000 * (1 + r) ^ 18
divide both sides of the equation by 250,000 to get:
4 = (1 + r) ^ 18
take the 18th root of both sides of the equation to get:
4 ^ (1/18) = (1 + r)
solve for (1 + r) to get:
1 + r = 1.080059739
that's the annual growth rate.

confirm by replacing (1 + r) in the equation with that to get:
f = 250,000 * 1.080059739 ^ 18 = 1,000,000
this confirms the value of (1 + 4) is correct.

1 + r is the annual growth factor.
subtract 1 from that to get the annual growth rte.
that would be .080059739.
multiply that by 100 and round to decimal places to get 8.01% growth rate per year.